1) Foreach of the production functions below, draw the isoquant passing througb the point z^(4,1). Label at least two points on the isoquant. Also determine whether the technology exhibits CRS,IRS or DRS. a. f(x)- 2x2 b. f(x)-x1/2+X2 c. f(x)- max(xiX2) d. f(x)-xiX22 2) Eoreach of the production functions below, find the cost function and conditional factor demands if w1-2 and w2-4. What is the amount of x1 and x2 that minimizes the cost of producing 4 units of output? a. f(x)-xiX2 b. f(x)- 2xı+X2 c. f(x)-min(xi,2x2) d. u(x)- max(xi,X2)
Answer #1

Answer:

a)given

f(x)=2*X2

For Z=(4,1)

So f(z)=f(4,1)=2*1=2

for Iso quant curve Q=2 units

So curve

Let the input is increased by '3' times then New inputs is 12,3  so new output f(12,3)=2*(3)=2*3=6 units

Since increase in output is also 3 times of earlier output which is equal to increase in inputs so This is constant return to scale(CRS).

B)

given

f(x)=X1^(1/2)+ X2

For Z=(4,1)

So f(z)=f(4,1)=4^(1/2)+1=3 units

for Iso quant curve Q=3 units

So curve

Iso Quant Curve 2.5 1.5 0.5 X1

Let the input is increased by '3' times then New inputs is (12,3) so new output f(12,3)=12^(1/2)+3=6.46 units

Since increase in output is only 6.46/3=2.15  times of earlier output which is less than  increase in inputs so This is Decreasing return to scale(DRS).

C)

given

f(x)=max(X1,X2)

For Z=(4,1)

So f(z)=f(4,1)=Max(4,1)=4

for Iso quant curve Q=4 units

So Curve is

Iso Quant Curve 4.5 3.5 2.5 1.5 0.5 X1

Let the input is increased by '3' times then New inputs is 12,3  so new output f(12,3)=Max(12,3)=12 units

Since increase in output is also 3 times of earlier output which is equal to increase in inputs so This is constant return to scale(CRS).

D)

given

f(x)=X1*X2^2

For Z=(4,1)

So f(z)=f(4,1)=4*1^2=4 units

for Iso quant curve Q=4 units

So Curve is

Let the input is increased by '3' times then New inputs is 12,3  so new output f(12,3)=12*3^0.5=20.78 units

Since increase in output is only 20.78/3=6.92  times of earlier output which is greater than increase in inputs so This is increasing return to scale(CRS).